An infrared heater has a surface area of 0.06m², and the emissivity of the surface is 0.85. how hot must the surface be for the heater to radiate with a power of 500W


1
Expert's answer
2021-05-30T13:22:41-0400

According to Stephen-Boltzmann law:


"q=\sigma A\epsilon(T_2^4-T_1^4)"

If the surrounding temperature is 0°C, or 273 K, the temperature of the heater must be


"T_2=\sqrt[4]{\frac{q}{\sigma A\epsilon}+T_1^4}=650\text{ K},"

or 377°C.